OpenStudy (anonymous):

there are 30 students in a class, the ratio for boys and girls are 3:8, how many girls are there??

OpenStudy (dumbcow):

8/11 = x/30 cross multiply

myininaya (myininaya):

or... b=boys g=girls b+g=30 and b/g=3/8 assuming its boys to girls and not girls to boys you have two equations and two unknowns have fun!

OpenStudy (anonymous):

what? this makes no sense lol there are decimals in the answers! haha

myininaya (myininaya):

i know i see this you have 240/11 girls lol

OpenStudy (anonymous):

don't you think something is wrong with the data in the question? 30 is NOT divisible by 11

OpenStudy (anonymous):

all i know is i wish my classes had a boy to girl ratio of 3:8 *wink*

OpenStudy (anonymous):

correct the question to something like "the ratio of boys and girls IS 3:7" then you would get an answer.

OpenStudy (anonymous):

240/11 girls....1 girl is a midget...she counts as 0.818 of a girl...therefore, 21 normal girls, 1 midget girl, 8 boys....problem solved. =D

myininaya (myininaya):

there would be a midget boy who counts .1818

myininaya (myininaya):

girls are worthed more than boys lol